OpenStudy (anonymous):

Explain how you would solve the problem below: Nathan has $16 more than Jeff. Jeff has $10 more than Brad. Together the Boys have $66. How much money does Brad have?

OpenStudy (anonymous):

This is the third time im asking this question, and i still dont understand

OpenStudy (anonymous):

well start with N = J + 16 then J = B+10 then B + J + N = 66 ( the letters are the first names of the boys). Now use a powerful mathematical tool known as subsititution!

OpenStudy (anonymous):

what is that. thats the thing i never got

OpenStudy (anonymous):

okay look at N = J + 16 , we know that J = B+10 right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so in N = J + 16, replace J. we know J = B+10 so we just put that were J is. N = (B+10) +16

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

so now we know N = B +10+16 and we know J = B+10 so go back to J+B+N = 66 and put in what J and N equal in terms of B, solve for B

OpenStudy (anonymous):

:OOMG I ACTUALLY GOT IT :DD THANK YOUUUU

OpenStudy (anonymous):

the more you practice the easier it becomes and then it'll be just as easy as adding 1+1